# Model Diffusion for Certifiable Few-shot Transfer Learning

### v1

@@ 4 Related Work
Risk certificates. Developing risk certificates for prediction models has been a central topic in theoretical machine learning (Vapnik, 1995; Shalev-Shwartz & Ben-David, 2014; Mohri et al., 2018) where rigorous theorems such as VC, Rademacher, and PAC-Bayes bounds have been proposed and studied. Traditionally these have been applied to simple models (e.g., linear classifiers or decision trees), however, there have been recent attempts to extend these theorems to large-scale deep models for non-vacuous risk certificates. For instance, in  Perez-Ortiz et al. (2021) the high complexity of KL divergence on large deep models has been circumvented by introducing a data-dependent prior that is learned by a held-out portion of training data. Recently  Lotfi et al. (2024) showed that parameter-level quantization of PEFT adapters can lead to non-vacuous bounds using the Kolmogorov complexity bounds. However, most of these approaches still require a large amount of training data to attain non-vacuous bounds. They fail in the low data regime considered in our paper. Within multi-task transfer learning scenarios similar to ours, a meta learning extension of the PAC-Bayes bound was proposed in Zakerinia et al. (2024). However, their method was only demonstrated for small-scale toy problems, and it easily incurs out-of-memory computational issue due to the nested gradient computations.
Model diffusion methods. Recently there have been several attempts to learn a diffusion model for generating model parameters. The neural net diffusion (Wang et al., 2024) trained a diffusion model for BN modules of ResNet-18 where the model parameter data are collected in the course of the SGD training, hence not very diverse. The ProtoDiff (Du et al., 2023) aimed for few-shot classification where based on ProtoNet (Snell et al., 2017) they aimed to diffuse the prototype vectors. As such the method is highly tied to and geared towards the ProtoNet few-shot learning, and hard to extensible to general model architectures. In MetaDiff (Zhang et al., 2024) they aimed for avoiding the Hessian computation in meta learning by treating the inner-loop gradient descent process as a reverse diffusion process. With the conditional sampling, they approached the meta learning problem. But to the best of our knowledge, none of the previous works offer risk certificates for the generated models from the diffusion as we did in our work.
Sparse adapter (PEFT) methods. Sparse adapter models aka PEFT adapters are widely used in practical few-shot adaptation, and effectively reduce the learnable parameters of neural networks. They are also an important component of our approach because learnable parameters should be small enough to be sampled by a diffusion model. In LoRA (Hu et al., 2021) left/right low rank matrices are trainable parameters. In VeRA (Kopiczko et al., 2024), they rather randomly choose and fix these two matrices while training only the diagonal matrices. This will greatly save the size of trainable parameters from LoRA. But still too large to be used as diffusion model state space. We adopt an even more sparse adapter model LoRA-XS (Balazy et al., 2024) which performs the singular value decomposition, replaces the diagonal singular value matrix with a trainable full matrix. Such adapters have facilitated large-data guarantees (Lotfi et al., 2024), but in our framework they will facilitate low-shot guarantees.
@@ 5 Experiments
We defer all architectural details, and hyperparameters to Appendix B.
@@ 5.1 Guarantees for few-shot LLM adaptation
Figure 2: Distribution of generalisation guarantees (x-axis, log scale) obtained over few-shot LLM adaptation episodes. Vertical lines indicate the vacuous bound threshold. Our STEEL learner provides a dramatically better distribution of outcomes in terms of provable generalisation compared to alternatives.
Table 1: Results on the LaMP LLM adaptation benchmark.
@@
Datasets: For low-shot LLM adaptation, we use the LaMP personalization benchmark (Salemi et al., 2024). LaMP contains a fixed number of users per dataset that is split over training (seen) and evaluation (unseen) clients. Training and evaluation clients are mutually exclusive. Each client has their own support data and query data. We choose three datasets from the benchmark, namely LaMP-2 (Personalized Movie Tagging), LaMP-3 (Personalized Product Rating), and LaMP-5 (Personalized Scholarly Title Generation). These make nominal classification, ordinal classification, and text generation tasks respectively.
@@
Setup: Following recent work on LaMP (Tan et al., 2024; Salemi et al., 2024), we first build a base model with task-specific capabilities by end-to-end fine-tuning of Flan-T5 base (Chung et al., 2024) on the user support data seen. Subsequently, to personalize the base model for a user, we train one LoRA-XS module on the support data of each training client (Bałazy et al., 2024). LoRA-XS rank of 6 and an alpha of 16 was used, producing a total of 2592 tunable parameters. We build the model zoo by collecting the LoRA-XS modules trained using seen users support data; same data that was aggregated to finetune the base model.
Baselines: We compare our proposal of using the model zoo directly, and our diffusion samples against two gradient-free methods, namely LoRA-Hub (Huang et al., 2024), MeZO (Malladi et al., 2023), and include SGD baseline for completeness. Please note that LoRA-Hub randomly samples from the model zoo and learns a new adapter as a linear combination of the sampled adapters. MeZO and SGD do not use the zoo. We refer the reader to Appendix B.2 for detailed hyperparameters. STEEL uses the bound described in Equation 3. Meanwhile, MeZO, SGD and LoRA-Hub use the quantization bound from Equation 5. Finally, for computational efficiency, we use Hierarchical search as described in Section 3.4.
@@
Results: Our main contribution relates to the ability to provably certify the generalisation of low-shot learning. In terms of low-shot LLM adaptation, Figure 2 visualises the distribution of certification outcomes over a large number of episodes for the three LAMP benchmarks. Taking note of the log-scale on the x-axis for generalisation guarantee strength, we can see that our STEEL learner provides dramatically better guarantees than conventional continuous-parameter learner alternatives, thanks to its discrete hypothesis space. The vertical lines indicate the threshold for vacuous bounds. Standard learners such as SGD and MeZO have no mass left of the threshold, while a substantial number of STEEL learning episodes are non-vacuously guaranteed. Table 1 provides more detailed quantitative results in terms of various metrics. To assess provable generalisation, the data visualised in Figure 2 is summarised as the % non-vacuous metric (the fraction of episodes which have guarantees with a strength above chance-level), and the median guarantee strength across episodes. To asses practical performance we report relevant metrics for each task such as Accuracy, RMSE, and ROGUE score.
@@
From these results we can see that: (1) Standard supervised learning approaches such as (gradient-based) SGD and (gradient-free) MeZO have no non-vacuous episodes - no few-shot learning task can be guaranteed. (2) Our STEEL model has the most non-vacuous episodes for each benchmark, with almost every few-shot learning episode being guaranteed in the LAMP-5 benchmarks. And the median STEEL episode has a substantially non-vacuous guarantee for all three benchmarks. (3) Interesingly, LoraHUB combined with Lotfi et al. (2024)’s discretization bound also has some non-vacuous episodes, but less than STEEL. (3) STEEL has comparable or better empirical test accuracy compared to existing approaches such as SGD and MeZO, while providing a huge improvement in certifiability.
@@ 5.2 Guarantees for few-shot visual recognition
Figure 4: “Learning curves” illustrating empirical and certified learning dynamics of STEEL with respect to samples/iterations, which is equivalent to hypothesis space size. More samples improves the training (support) error, while increasing the complexity penalty. The sum of these two terms instantiates the generalisation guarantee (Eq. 3) achieved for a given number of samples.
Method SGD BBPT STEEL
@@
Non-Vacuous Ratio 0.00% 0.00% 100.00%
Average Gap 2.49 2.48 0.24
Min Bound 2.55 2.55 0.30
Median Bound 2.58 2.59 0.36
Max Bound 2.64 2.68 0.47
Average Accuracy 90.32 89.27 88.40
@@
Non-Vacuous Ratio 0.00% 0.00% 100.00%
Average Gap 2.43 2.46 0.24
Min Bound 2.55 2.55 0.36
Median Bound 2.55 2.63 0.42
Max Bound 2.58 2.78 0.53
Average Accuracy 87.95 83.20 81.50
@@
Non-Vacuous Ratio 0.00% 0.00% 97.50%
Average Gap 2.31 2.45 0.22
Min Bound 2.58 2.64 0.45
Median Bound 2.64 2.80 0.61
Max Bound 2.78 3.01 0.85
Average Accuracy 65.57 62.37 61.37
@@
Non-Vacuous Ratio 0.00% 0.00% 100.00%
Average Gap 2.51 2.50 0.27
Min Bound 2.55 2.55 0.31
Median Bound 2.55 2.59 0.40
Max Bound 2.55 2.70 0.61
Average Accuracy 95.90 90.15 84.90
Table 2: CLIP+CoOp-based few-shot learning. Aggregates over multiple 16-Shots 5-way learning episodes.
Datasets: For vision, we use fine-grained classification datasets that have readily available training (seen) and novel (unseen) classes split. We choose CUBirds (Wah et al., 2011), FGVC-Aircraft (Maji et al., 2013), Describable Textures (Cimpoi et al., 2014) and Flowers-101 (Nilsback & Zisserman, 2008). We use the split offered by learn2learn for the first three datasets (Arnold et al., 2020), and the Flowers-101 split from Meta-Dataset (Triantafillou et al., 2020).
Setup: We sample random few-shot $n$-shot $k$-way learning tasks as per meta-learning literature (Triantafillou et al., 2020). We randomly select 5 classes (5-way) from a given split and for each class we randomly sample $n$-shots. We evaluate 1, 2, 4, 8 and 16-shots. The model zoo is built by sampling from the training classes of the aforementioned datasets. Meanwhile, we evaluate on unseen classes and samples. Please note that training and evaluation classes are mutually exclusive. For a given task, we use CoOp for adaptation (Zhou et al., 2022) which is simply prompt tuning for CLIP. We finetune a 2-token prompt appended in front of the class name for every task to build a model zoo. This results in total tunable parameters of length 1024. For vision experiments, we found that the forward passes were fast enough to conduct exhaustive search over sampled prompts from the diffusion model.
Baselines: We compare our proposal of using the model zoo directly, and our diffusion samples against Black-Box Prompt Tuning (BBPT) (Yu et al., 2023) which is a gradient-free version of CoOp, optimizing the prompt using evolutionary optimization. We also include SGD based CoOp for completeness. For details on methods hyperparameters, we refer the reader to Appendix B.3. STEELuses the bound described in Equation 3. Meanwhile, BBPT and SGD use the quantization bound from Equation 5.
Results: The results in terms of mean training and testing error versus complexity are summarised for three datasets in Figure 1. The dots for each learner reflect the training set sizes of 1, 2, 4, 8, and 16-examples per-class, and the white/grey zone separation delineates the space of non-vacuous vs vacuous bound outcomes. The main message is that only our finite hypothesis class approaches achieve any non-vacuous guarantees across this whole range of training set sizes. Every result for the standard SGD and BBPT approaches is vaccuous and cannot be guaranteed.
For the 16-shot case, these experiments are quantified in more detail in Table 2. Similarly to the results for LLM adaptation, we can see the dramatic difference in % of non-vacuous episodes, and dramatic improvement in the min, median and max bound obtained over episodes. Compared to the LLM case, our STEEL pays a slighly higher price here in terms of empirical test accuracy compared to SGD for some benchmarks, however this is still small compared to the stark difference in certification performance.
@@
Further Analysis: We finally discuss and provide some insight into the learning process of our discrete hypothesis class learner. Standard gradient-descent takes repeated update steps to find a model that better fits a training set. By analogy, our STEEL gradient-free learner would draw more samples as it attempts to iteratively sample a model that better fits the training set. Our main experiments use a fixed number of 20,000 samples on all vision datasets, but Figure 4 illustrates our learner’s behaviour by showing the equivalent of a learning curve for our model. The x-axis is the number of samples drawn, and equivalently the learning theoretic hypothesis space size. Unlike SGD, this means that there is a direct dependence of hypothesis class complexity ($|\Theta|$ in Eq. 3) and the number of iterations/samples. This is reflected in the steadily increasing red complexity curve in Figure 4(left, middle). We can also see that the training/support error goes down consistently over iterations/samples as the sampler progressively discovers better models. The generalisation bound (black line) is given by the sum of the training error and complexity. The figure illustrates one case (Flowers, middle, right) where the bound continues to improve up to a large number of samples/hypothesis size, because the continued improvement in training error outweighs the complexity gain. It also illustrates a case (DTD, left) where the training error improvement is slower and quite rapidly outweighed by the complexity gain, so that the best bound is actually achieved after quite a small number of samples.
@@
Sampler vs Zoo: Our approach compresses the upstream set of pre-trained models into a learned model generator. Selecting among the upstream models using downstream task performance as a criterion provides an alternative approach to learning that also corresponds to a finite hypothesis space. Our generator approach was motivated by ensuring scalability with respect to a large number of upstream models, and also to improve accuracy by enabling interpolation between upstream models rather than solely being limited to selecting one of them. Figure 1 shows that STEEL’s diffusion sampler tends to provide improved accuracy compared to its raw model zoo, especially for Flowers and DTD. On average across all datasets STEEL consistently outperforms Model Zoo. Detailed per-dataset per-shot performance is deferred to Appendix C, Table 3.

### v2

@@ 4 Related Work
Risk certificates. While traditionally applied to simple models, recent work extends risk certificates to deep learning. Notable approaches include using data-dependent priors [21] and parameter quantization of PEFT adapters [15] in tabula-rasa and single source transfer learning respectively. Nevertheless, these still require large training sets to obtain non-vaucous bounds. In the meta-learning literature,  Zakerinia et al. [39] proposed meta-learning PAC-Bayes bounds, which aims to facilitate tighter generalisation bounds by extracting a common prior from up-stream tasks in a multi-task setting. However, it was only demonstrated on toy problems and doesn’t scale to large models due to the memory requirements of nested gradient computations.
Model diffusion methods. Several works explore diffusion for generating model parameters: NNDiffusion [35] for BN modules, ProtoDiff [6] for ProtoNet-based few-shot learning, and MetaDiff [40] and D2NWG [28] for gradient-free meta-learning. Scaling properties of Diffusion based learning were also explored [25]. However, none provide risk certificates for the generated models.
Sparse adapter (PEFT) methods. Sparse adapters reduce learnable parameters, crucial for our diffusion-based sampling. While LoRA [11] uses trainable low-rank matrices and VeRA [13] uses fixed matrices with trainable diagonals, we adopt LoRA-XS [3], which uses SVD with a trainable full matrix for the singular values. Such adapters have facilitated large-data guarantees [15], but in our framework they will facilitate low-shot guarantees.
Connection to our Approach. In terms of assumptions, we consider the same multi-task setting of meta-learning [10], model-zoo [12] and model diffusion [35] approaches. They all aim to facilitate downstream learning by knowledge transfer from upstream tasks. In terms of solution we share the benefit of model-zoo and model-diffusion methods in being able to use third-party off-the-shelf pre-trained upstream models, rather than requiring an expensive joint learning procedure like meta-learners. Like meta-learners and model-diffusers, but unlike zoo methods, we do learn a task-agnostic component (the diffusion model).
@@ 5 Experiments
Metrics for Certified Learning Quality: In the low-shot learning context, we run many learning episodes with different random small training sets [36]. Thus we need evaluation metrics for the typical empirical performance and certificate strength instead of a single accuracy and certificate for one large scale learning [15]. This is straightforward for empirical test performance: we report the average over episodes of the relevant task metric (e.g., accuracy, RMSE). For certificates, we report the minimum (best), median (typical) and maximum (worst) case error guarantee over all downstream tasks. Since many learners often produce vacuous certificates in the low-shot regime (e.g., guaranteed error not below 1), we use the proportion of tasks which have a non-vacuous certificate as our leading metric. Finally, to quantify how tight the certificates are, we also report gap - the average distance between certificates and query errors over all faithful bounds. More concretely, this is computed as $\sum_{i=0}^{K}\frac{\text{max}(0,\text{bound}_{i}-\text{query error}_{i})}{K}$ where $K$ is the total number of downstream tasks.
Competitors: For single task methods that exploit only the foundation model and the target downstream task, we compare standard SGD, along with MeZO [17] and BBPT [38] – two state of the art gradient-free learners – for fine-tuning. For multi-task alternatives that exploit also the same task distribution as STEEL, we compare the model zoo learner LoRA-Hub [12] and the meta-learner MetaPB [39]. LoRA-Hub randomly samples from the model zoo and learns a new adapter as a linear combination of sampled adapters. Note that all methods are already combined with the same (experiment-specific) PEFT strategy for fair comparison. For SGD, MeZO, BBPT and MetaPB, this is an upgrade provided by us - they would definitely otherwise fail in the full parameter space. For SGD, MeZO, BBPT and LoRA-Hub, we combine them with recent quantization bound of [15] (Eq. 7) to provide guarantees, MetaPB uses their own PAC Bayes bound, and STEEL uses the bound in Eq. 3.
Figure 2: Distribution of generalisation guarantees (x-axis, log scale) obtained over few-shot LLM adaptation episodes. Vertical lines indicate the vacuous bound threshold. STEEL provides a dramatically better distribution of provable generalisation outcomes compared to alternatives.
@@ 5.1 Guarantees for few-shot LLM adaptation
Datasets: For low-shot LLM adaptation, we use the LaMP personalization benchmark [24]. LaMP contains a fixed number of users per dataset that is split over training (seen) and evaluation (unseen) clients. Training and evaluation clients are mutually exclusive. Each client has their own support data and query data.
Table 1: LaMP LLM adaptation benchmark results.
@@
We choose three datasets from the benchmark, namely LaMP-2 (Personalized Movie Tagging), LaMP-3 (Personalized Product Rating), and LaMP-5 (Personalized Scholarly Title Generation). These make nominal classification, ordinal classification, and text generation tasks respectively.
@@
Setup: Following recent work on LaMP [30, 24], we first build a base model with task-specific capabilities by end-to-end fine-tuning of Flan-T5 base [4] on the user support data seen. Subsequently, to personalize the base model for a user, we train one LoRA-XS module on the support data of each training client [3]. LoRA-XS rank of 6 and an alpha of 16 was used, producing a total of 2592 tunable parameters. We build the model zoo by collecting the LoRA-XS modules trained using seen users support data; same data that was aggregated to finetune the base model. For computational efficiency, STEEL uses Hierarchical search as described in Section 3.4. We refer the reader to Appendix B.2 for detailed hyperparameters.
@@
Results: Our main contribution relates to the ability to provably certify the generalisation of low-shot learning. In terms of low-shot LLM adaptation, Figure 2 visualises the distribution of certification outcomes over a large number of episodes for the three LAMP benchmarks. Taking note of the log-scale on the x-axis for generalisation guarantee strength, we can see that our STEEL learner provides dramatically better guarantees than conventional continuous-parameter learner alternatives, thanks to its discrete hypothesis space. The vertical lines indicate the threshold for vacuous bounds. Standard learners such as SGD and MeZO have no mass left of the threshold, while a substantial number of STEEL learning episodes are non-vacuously guaranteed. Table 1 provides more detailed quantitative results in terms of various metrics. Notably, to assess provable generalisation, the data visualised in Figure 2 is summarised as the % non-vacuous metric (the fraction of episodes which have guarantees with a strength above chance-level), and the median guarantee strength across episodes.
@@
From Table 1 results we can see that: (1) Standard supervised learning approaches such as (gradient-based) SGD and (gradient-free) MeZO have no non-vacuous episodes - no few-shot learning task can be guaranteed. (2) STEEL has the most non-vacuous episodes for each benchmark, with almost every few-shot learning episode being guaranteed in the LAMP-5 benchmarks. The median STEEL episode also has a substantially non-vacuous guarantee for all three benchmarks. (3) Interestingly, LoraHUB combined with Lotfi et al. [15]’s discretization bound also has some non-vacuous episodes, but less than STEEL. (3) STEEL has comparable or better empirical test accuracy compared to existing approaches such as SGD and MeZO, while providing a huge improvement in certifiability.
@@ 5.2 Guarantees for few-shot visual recognition
Figure 4: “Learning curves" illustrating empirical and certified learning dynamics of STEEL with respect to samples/iterations, which is equivalent to hypothesis space size. More samples improves the training (support) error, while increasing the complexity penalty. The sum of these two terms instantiates the generalisation guarantee (Eq. 3) achieved for a given number of samples.
Datasets: For vision, we use fine-grained classification datasets that have readily available training (seen) and novel (unseen) classes split. We choose CUBirds [34], FGVC-Aircraft [16], Describable Textures [5] and Flowers-101 [20]. We use the split offered by learn2learn for the first three datasets [2], and the Flowers-101 split from Meta-Dataset [31].
Setup: We sample random few-shot $n$-shot $k$-way learning tasks as per meta-learning literature [31]. We use CLIP as our foundation model [22]. We randomly select 5 classes (5-way) from a given split and for each class we randomly sample $n$-shots. We evaluate 1, 2, 4, 8 and 16-shots. The model zoo is built by sampling from the training classes of the aforementioned datasets. Meanwhile, we evaluate on unseen classes and samples. Please note that training and evaluation classes are mutually exclusive. For a given task, we use CoOp PEFT adaptation [41] which is simply prompt tuning for CLIP. We finetune a 2-token prompt appended in front of the class name for every task to build a model zoo. This results in total tunable parameters of length 1024. For vision experiments, we found that the forward passes were fast enough to conduct exhaustive search over sampled prompts from the diffusion model. For Meta-PB, we detail the meta-training setup, downstream adaptation and Meta-PB bounds in Appendix A.3. For details on all methods hyperparameters, we refer the reader to Appendix B.3.
Results: The results in terms of mean training and testing error versus complexity are summarised for three datasets in Figure 1. The dots for each learner reflect the training set sizes of 1, 2, 4, 8, and 16-examples per-class, and the white/grey zone separation delineates the space of non-vacuous vs vacuous bound outcomes. The main message is that only our finite hypothesis class approaches achieve any non-vacuous guarantees across this whole range of training set sizes. Every result for the standard SGD and BBPT approaches is vaccuous and cannot be guaranteed. Please note, Meta-PB is not included as it is not directly comparable on the plot y-axis (empirical error term); Meta-PB uses the unbounded cross-entropy loss (Equation 23 as the risk metric, while our and other competing methods use empirical error.
For the 16-shot case, these experiments are quantified in Table 2. Similarly to the results for LLM adaptation, we can see the dramatic difference in % of non-vacuous episodes, and dramatic improvement in the min, median and max bound obtained over episodes. Compared to the LLM case, STEEL pays a slighly higher price in terms of empirical test accuracy compared to SGD for some benchmarks, however this is small compared to the stark difference in certification performance. The state of the art Meta-PB, also oriented at low-shot certification in some cases (CUB) manages to non-vacuously certify most episodes. However, in all cases the strength of this certification is much worse than STEEL (gap, median bound).
Method SGD BBPT Meta-PB STEEL
@@
Non-Vacuous Ratio 0.00% 0.00% 97.50% 100.00%
Average Gap 2.49 2.48 0.45 0.24
Min Bound 2.55 2.55 0.34 0.30
Median Bound 2.58 2.59 0.55 0.36
Max Bound 2.64 2.68 0.85 0.47
Average Accuracy 90.32% 89.27% 89.24% 88.40%
@@
Non-Vacuous Ratio 0.00% 0.00% 50.00% 100.00%
Average Gap 2.43 2.46 0.54 0.24
Min Bound 2.55 2.55 0.57 0.36
Median Bound 2.55 2.63 0.80 0.42
Max Bound 2.58 2.78 1.06 0.53
Average Accuracy 87.95% 83.20% 81.12% 81.50%
@@
Non-Vacuous Ratio 0.00% 0.00% 0.00% 97.50%
Average Gap 2.31 2.45 0.86 0.22
Min Bound 2.58 2.64 0.87 0.45
Median Bound 2.64 2.80 1.25 0.61
Max Bound 2.78 3.01 1.76 0.85
Average Accuracy 65.57% 62.37% 59.84% 61.37%
@@
Non-Vacuous Ratio 0.00% 0.00% 10.00% 100.00%
Average Gap 2.51 2.50 0.68 0.27
Min Bound 2.55 2.55 0.65 0.31
Median Bound 2.55 2.59 1.15 0.40
Max Bound 2.55 2.70 1.77 0.61
Average Accuracy 95.90% 90.15% 71.23% 84.90%
Table 2: Few-shot visual recognition. Aggregate over 16-Shots 5-way learning episodes.
@@ 5.3 Further Analysis
Learning Curves: We discuss and provide some insight into the learning process of our discrete hypothesis class learner. Standard gradient-descent takes repeated update steps to find a model that better fits a training set. By analogy, our gradient-free STEEL draws more samples as it attempts to iteratively sample a model that better fits the training set. Our main experiments use a fixed number of 20,000 samples on all vision datasets. Figure 4 illustrates our learner’s behaviour by showing the equivalent of a learning curve for our model. The x-axis is the number of samples drawn, and equivalently the learning theoretic hypothesis space size. Unlike SGD, this means that there is a direct dependence of hypothesis class complexity ($|\Theta|$ in Eq. 3) and the number of iterations/samples. This is reflected in the steadily increasing red complexity curve in Figure 4(left, middle). We can also see that the training/support error goes down consistently over iterations/samples as the sampler progressively discovers better models. The generalisation bound (black line) is given by the sum of the training error and complexity. The figure illustrates one case (Flowers, middle, right) where the bound continues to improve up to a large number of samples/hypothesis size, because the continued improvement in training error outweighs the complexity gain. It also illustrates a case (DTD, left) where the training error improvement is slower and quite rapidly outweighed by the complexity gain, so that the best bound is actually achieved after quite a small number of samples.
@@
Sampler vs Zoo: Our approach compresses the upstream set of pre-trained models into a learned model generator. Selecting among the upstream models using downstream task performance as a criterion provides an alternative approach to learning that also corresponds to a finite hypothesis space. Our generator approach was motivated by ensuring scalability with respect to a large number of upstream models, and also to improve accuracy by enabling interpolation between upstream models rather than solely being limited to selecting one of them. Figure 1 shows that STEEL’s diffusion sampler tends to provide improved accuracy compared to its raw model zoo, especially for Flowers and DTD. On average across all datasets STEEL consistently outperforms Model Zoo. Detailed per-dataset per-shot performance is deferred to Appendix C, Table 3. Furthermore, we show in Appendix C.2 how to certify zero-shot CLIP using confidence intervals, and demonstrate that STEEL not only achieves tighter guarantees but also substantially outperforms it in terms of empirical performance.
Figure 5: Support/query error and certified risk vs. support set size on iNaturalist birds. As the number of shots increases, support and query errors converge, while the bound continues to tighten. The gap between the certified risk and empirical query error drops to 6% at 128 shots, demonstrating the ability to produce tight certificates.
Tightening STEEL Certificates: We focused thus far on the very low-shot regime to emphasise the unique achievement of obtaining the first non-vacuous certificate in this extreme setting. Nevertheless, stronger certificates may be necessary in many applications. We therefore explore increasing the training data from very low to moderately low shot. Specifically, we evaluate the iNaturalist dataset [9] (birds subset, with overlapping classes excluded) with a diffusion model pre-trained on CUB, which facilitates increasing shots to 128 examples. Results in Figure 5 show that: (1) As the training set grows, the support error increases initially as there are more data points to fit, but soon stabilizes. The finite hypothesis-space does not problematically under-fit an increasingly large training set. (2) Query error follows a similar trend, decreasing with support set size and saturating around 64 shots. (3) Crucially, the STEEL bound continues to improve with more shots. The increasing training set size continues to shrink the certificate’s complexity term, even after the support error has saturated, thus leading to continued tightening of the bound. At 64 or 128 shots, the certified risk is tight – only 6% larger than the empirical test error, while still using far fewer examples than standard certification approaches [15]. (4) This good result is obtained despite the distribution-shift (CUB vs iNaturalist) between upstream and downstream data. STEEL is robust to to imperfect upstream-downstream task distribution alignment.